Circular No. 8383

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 P1 (NEAT) K. J. Lawrence, Jet Propulsion Laboratory, reports the discovery by

the NEAT project of a slightly diffuse comet (coma diameter ∼ 3′′; discovery observation below) on CCD images taken with the Palomar 1.2-m Schmidt telescope. Following posting on the ‘NEO Confirmation Page’, other CCD observers have remarked on the object’s cometary appearance, including J. Tichá and M. Tichý at Kleť (1.06-m telescope; object diffuse with an asymmetric coma or faint tail toward p.a. 250^{o} and red mag 18.7–18.8 on Aug. 7.92 UT; coma diameter 13′′ and faint tail in p.a. 250^{o} on Aug. 8.88); P. Birtwhistle at Great Shefford, Berkshire, England (0.30-m f/6.3 reflector; object distinctly diffuse with a diameter of 10′′ — three times the size of star images of similar brightness on co-added images taken on Aug. 7.94– 7.96); and P. Kušnirák at Ondřejov (0.65-m f/3.6 reflector; diffuse with coma diameter 10′′ on Aug. 7.93 and 7.97).

2004 UT  α_{2000}  δ_{2000}  Mag.
Aug. 5.32080  21^{h}42^{m}35.^{s}46  +12^{o}44′16′′.6  19.9

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-P23.

T = 2006 May 13.024 TT  ω = 99.^{o}711
Ω = 290.544 }2000.0
q = 5.69231 AU  i = 32.213

SUPERNOVA 2004dp T. Puckett and M. Peoples, Mountain Town, GA, report the discovery

of an apparent supernova (mag 18.2) on unfiltered CCD frames taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Aug. 7.21 and 8.19 UT. The new object is located at α = 21^{h}33^{m}52^{s}.27, δ = +41^{o}51′01′′.4 (equinox 2000.0), which is 3′′.7 west and 6′′.3 south of the center of the apparent host galaxy. SN 2004dp was not present on nine images taken by Puckett from 2000 to 2003 (limiting mag ∼ 20.0).

SUPERNOVA 2004dg IN NGC 5806 Additional unfiltered CCD magnitudes of SN 2004dg (cf. IAUC 8375):

June 14.673 UT, [18.3 (Itagaki); July 22.493, 17.4 (Itagaki); Aug. 7.844, 17.6 (E. Prosperi, Larciano, Italy).

2004 August 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green